# VoxEval: Benchmarking the Knowledge Understanding Capabilities of End-to-End Spoken Language Models

### v3

@@ Abstract
Wenqian Cui1, Xiaoqi Jiao2, Ziqiao Meng3, Irwin King1, 1The Chinese University of Hong Kong, 2LIGHTSPEED STUDIOS, 3National University of Singapore
@@ 1 Introduction
Existing benchmarks Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024) for evaluating audio-based knowledge comprehension face two major limitations. First, they fail to support end-to-end speech evaluations. These benchmarks typically build datasets by matching audio clips with text-based question/answer pairs, which differ from the speech interaction pipeline. Second, they do not account for diverse input audio conditions, such as variations in speaker timbre, speaking styles, and audio quality, which are common in real-world settings. However, it is crucial for SLMs to demonstrate robustness against these varying conditions, producing consistent answers when the input content remains the same Hurst et al. (2024). Additionally, we identify a unique challenge in SLM knowledge evaluation: not all knowledge comprehension topics are straightforward for SLM evaluation. For example, mathematical reasoning is easy to evaluate in written form due to the concise nature of mathematical expressions. However, evaluating it in spoken form remains a challenge. Despite this, possessing math reasoning skills is crucial for SLMs, as they are essential for applications like tutoring and simple problem-solving. Figure 1 illustrates the three key elements mentioned above.
Based on the key elements outlined, we introduce VoxEval, a SpeechQA benchmark designed to assess the knowledge understanding capabilities of SLMs with the following properties. First, VoxEval includes audio recordings for both questions and answers, enabling end-to-end speech interaction evaluation. Second, it incorporates diverse input audio conditions for the same question, allowing us to test the robustness of SLMs across various audio scenarios. Third, to tackle the challenges of evaluating subjects like mathematics and other complex topics, we identify the critical distinctions between written and spoken formats, transform written content into its spoken equivalent, and incorporate Chain-of-Thought prompts to assess the reasoning ability of SLMs in a step-by-step problem-solving setting. Table 1 presents the statistics for VoxEval, while Table 2 compares the distinctive features of VoxEval with those of other benchmarks. Our evaluation of recently developed SLMs shows that VoxEval poses a substantial challenge for current models, as they perform poorly on VoxEval, exhibit sensitivity to changes in input audio conditions, and demonstrate very limited reasoning capabilities. To summarize our contributions:
@@ 2.1 Speech Large Language Models
SLMs e2e # KUPs # variations math CoT
@@
Table 2: A comparison of our benchmark with other benchmarks. NUP and CoT stand for Knowledge Understanding Problem and Chain-of-Thought prompts.
@@ 2.2 SpeechLLM Evaluation Benchmarks
The evaluation of SpeechLLMs centers on their ability to model speech data, with researchers examining various aspects and levels. Initially, multiple datasets are utilized to assess the linguistic modeling capabilities of SpeechLLMs. For instance, sWUGGY Nguyen et al. (2020) is employed for lexical evaluation, sBLIMP Nguyen et al. (2020) for syntactical assessment, and sStoryCloze Hassid et al. (2024) for semantic evaluation. Additionally, the modeling of paralinguistic elements (spoken content beyond words) is crucial for SpeechLLMs, prompting the development of datasets that evaluate paralinguistic information from both token and perceptual levels Kharitonov et al. (2022); Nguyen et al. (2024). Beyond the modeling aspect, the performance of downstream applications is vital for the practical use of SpeechLLMs in real-world scenarios. Recently, numerous downstream evaluation benchmarks have emerged Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024), focusing on general audio understanding, including speech, environmental sounds, and music. However, as highlighted in Section 1, these benchmarks contain all QA pairs in text format, making them suitable for evaluating S2TLLMs but not SLMs. The closest benchmark to our work is VoiceBench Chen et al. (2024), which also evaluates end-to-end SLMs under various audio conditions. However, our study places a stronger emphasis on evaluating the knowledge comprehension abilities of SLMs, not only by including a broader range of topics but also by particularly focusing on their complex skills like mathematics. We also offer a detailed analysis of the performance of SLMs across various scenarios.
@@ 3.1 Data Construction
Building a speech-based knowledge understanding benchmark requires the collection of knowledge from different subjects and the construction of question-answer pairs about the knowledge in speech format. Given that there are numerous existing textual knowledge understanding benchmarks (see Section 2.3), it is logical to build the speech-based benchmark upon textual benchmarks. In this work, we leverage the MMLU Hendrycks et al. (2021) dataset to build VoxEval. Specifically, we transform the questions from MMLU into speech using the OpenAI Text-to-Speech (TTS) API  [^2] https://platform.openai.com/docs/guides/text-to-speech. We select MMLU as the underlying dataset for the following reasons:
@@
To transform MMLU into its spoken version, we need to concatenate all the text within a problem into a single sequence. In MMLU, every data is a Multiple Choice Question (MCQ) with four answer choices. We use a simple sentence to concatenate the question and answer choices. Therefore, every converted data looks like this:
@@ 3.2 Various Input Conditions
As mentioned in Section 1, it is crucial for the answers of SLMs to be robust under various input audio conditions. Inspired by this, we incorporate MMLU questions with different input audio conditions into VoxEval. Specifically, we consider the following types of input conditions.
@@ 3.2.2 Different Speaking Styles
In real-world conversations, even a single individual may use different speaking styles depending on the situation. Therefore, Speech Language Models (SLMs) need to manage various input styles. We identify two types of style variations: linguistic and paralinguistic.
@@
SLMs SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice
@@
Alloy 0.01 4.80 20.84 12.16 37.63
Echo 0.01 5.58 20.96 12.21 37.64
Fable 0.00 1.16 20.84 11.53 36.42
Nova 0.01 3.32 20.70 12.98 36.77
Onyx 0.02 2.75 19.66 11.92 37.64
Shimmer 0.00 5.16 20.76 12.05 38.15
@@
Linguistic 0.01 4.88 20.44 11.87 36.43
Speed 0.01 5.03 19.11 10.13 34.69
Pitch 0.00 5.44 17.88 6.09 33.45
@@
Noise 0.00 3.68 19.50 10.18 36.95
Other Env Acou 0.01 4.34 20.19 10.51 37.28
@@
SLMs $S\rightarrow S$ $S\rightarrow T$ $T\rightarrow S$ $T\rightarrow T$
@@
  • Tempo change. In real-life situations, people often vary their speaking speed, sometimes talking faster and other times slower. To mimic these variations in speaking pace, we apply tempo changes. For each question input, we randomly select a tempo-changing rate that ranges from half the original speed to twice as fast.
@@ 3.2.3 Different Audio Qualities
  • Background noise. Sometimes the input audio may be accompanied by background noise, which can affect the clarity and intelligibility of the spoken content. We consider different kinds of noise, including Gaussian noise, colored noise, background music, and short interruption noise. For each problem, a kind of noise is randomly selected and applied to the original audio.
  • Other environment acoustics. We consider speech recorded under various conditions, including aliasing, room impulse response, low pass filter, high pass filter, band pass filter, bitcrush, gain, clipping distortion, and seven band parameter EQ. Full explanations and transformation settings are presented in Appendix A. A random audio effect is applied to each question input.
@@
SLMs (MGS) SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice
@@
Table 7: Average accuracy (%) for various subject groups across different input audio conditions are presented. The scores are shown in the format “STEM/Social Science/Humanities/Others" subjects. MGD stands for Mean Grouped Score. The highest score in each scenario is highlighted in bold.
@@ 3.3 Handle Math Expressions and More
Mathematical problems often include numerous mathematical expressions, and current TTS systems struggle to accurately convert these expressions into speech, as they are typically trained only on textual words. To tackle this issue, we propose a two-step method for synthesizing math problems. First, we convert the written math problems into their spoken form, and then we use the TTS system to generate speech. Specifically, we create a set of few-shot examples and use GPT-4o to handle the conversion process. Our focus is on transforming Arabic numerals, units, operators, and brackets, as illustrated by the examples in Table 4. The complete prompt is provided in Figure 3 and 4.
@@
SLMs Direct LLM Direct LLM Direct LLM
@@
SPIRIT-LM 14.29 14.02 16.30 14.44 18.00 14.00
Moshi 12.17 10.32 14.44 11.48 8.00 10.00
GLM-4-Voice 27.25 28.04 14.44 15.93 22.00 22.00
@@
SPIRIT-LM 15.08 14.02 16.30 15.19 16.00 17.00
Moshi 11.11 11.11 13.33 11.48 15.00 14.00
GLM-4-Voice 27.25 27.78 14.44 15.19 22.00 22.00
@@ 4.1 Experimental Setups
Models. We evaluate VoxEval on five recently introduced SLMs: SpeechGPT Zhang et al. (2023), TWIST Hassid et al. (2024), SPIRIT-LM Nguyen et al. (2024), Moshi Défossez et al. (2024), and GLM-4-Voice Zeng et al. (2024). These models represent diverse training approaches within the SLM domain. Specifically, SpeechGPT and TWIST are pre-trained using unsupervised speech data. SPIRIT-LM and GLM-4-Voice are pre-trained on interleaved speech and text data. Moshi is pre-trained on aligned speech-text data. All the models are built upon TLM checkpoints. Furthermore, while the other models process audio data through a single stream, Moshi uniquely represents audio data using multiple streams.
@@ 4.2 Results
To answer RQ1, we evaluate five SLMs on VoxEval, as shown in Table 5. Results reveal that current SLMs perform poorly, with most failing to surpass random guessing. Since VoxEval questions have four answer choices, random guessing yields an expected score of 25%. However, only GLM-4-Voice exceeds this baseline, indicating that SLMs often struggle to follow instructions and select the correct answer. This underscores the challenge of enabling SLMs to “speak" accurate responses, making VoxEval a tough benchmark. A case study of SLMs’ spoken responses is in Appendix C. Notably, SpeechGPT’s performance is near 0%. This is likely due to our approach of directly performing speech-to-speech generation without utilizing the “chain-of-modality" mode described in Zhang et al. (2023). We choose not to use the “chain-of-modality" approach because it involves converting the input into text before generating a response, which we believe would significantly increase inference latency and is not an optimal strategy for end-to-end SLMs.
@@
To answer RQ4, we present the performance of SLMs across three math subjects in VoxEval, as shown in Table 8. The selected subjects—elementary math, high school math, and college math—are intended to represent varying levels of reasoning complexity. The math-solving abilities of SLMs are tested under two scenarios: direct answer, where the model outputs the answer directly, and Chain-of-Thought prompting (CoT) Wei et al. (2022), which uses annotated intermediate steps from three dev-set questions per subject as examples appended before the actual question. All the CoT examples and questions are in speech format. CoT evaluations are conducted in two settings: Cut, where the input audio is trimmed to the last 80 seconds (consistent with other experiments), and No cut, where the audio is left untrimmed to preserve longer CoT prompts. Two metrics are used for evaluation: string matching, the standard final answer extraction method, and LLM, where GPT-4o Hurst et al. (2024) is used to extract answers to complement string matching, as models often include lengthy reasoning steps that hinder accurate string-based extraction when applying CoT. Our observations include:
  • Existing SLMs perform similarly across different levels of math questions. Interestingly, in some cases, SLMs perform better on more challenging questions. For example, TWIST and GLM-4-Voice achieve the highest performance in college math. This counterintuitive fact suggests that existing SLMs are unlikely to possess basic reasoning abilities.
@@ 6 Limitations
The limitations of this work come from the scope and the evaluations. First, the VoxEval benchmark focuses exclusively on assessing the robustness of SLMs in terms of knowledge understanding, leaving out other critical aspects such as fairness, toxicity, hallucination, etc. These issues are particularly relevant in generative scenarios, where SLMs may exhibit biases, generate harmful content, or produce inaccurate information. However, these aspects fall outside the scope of VoxEval and require separate evaluation frameworks tailored to address them comprehensively. Second, the evaluation of the four modality combinations (speech-to-speech, speech-to-text, text-to-speech, and text-to-text) is not comprehensive enough, as it only includes one SLM (SPIRIT-LM). Expanding this evaluation to a broader range of models would enhance the reliability and generalizability of the findings.
Appendix A Explanation for Various Speech Recording Conditions

### v4

@@ Abstract
Wenqian Cui1thanks: Corresponding author: wenqian.cui@link.cuhk.edu.hk, Xiaoqi Jiao2, Ziqiao Meng3, Irwin King1, 1The Chinese University of Hong Kong, 2LIGHTSPEED STUDIOS, 3National University of Singapore
@@ 1 Introduction
Existing benchmarks Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024) for evaluating audio-based knowledge comprehension face two major limitations. First, they fail to support end-to-end speech evaluations. Current benchmarks construct datasets by either pairing audio clips with text-based QA pairs or providing speech-based questions while evaluating text-based responses. This approach fundamentally misaligns with real-world speech interaction pipelines, where users expect seamless speech-to-speech communication. Evaluating SLMs end-to-end is crucial, as multiple studies have documented substantial performance degradation in SLMs when transitioning from speech-to-text to speech-to-speech evaluations Chen et al. (2025); Long et al. (2025). Second, they do not account for diverse input audio conditions, such as variations in speaker timbre, speaking styles, and audio quality, which are common in real-world settings. However, it is crucial for SLMs to demonstrate robustness against these varying conditions, producing consistent answers when the input content remains the same Hurst et al. (2024), especially for fact-based questions. Additionally, we identify a unique challenge in SLM knowledge evaluation: not all knowledge comprehension topics are straightforward for SLM evaluation. For example, mathematical reasoning is easy to evaluate in written form due to the concise nature of mathematical expressions. However, evaluating it in spoken form remains a challenge. Despite this, possessing math reasoning skills is crucial for SLMs, as they are essential for applications like tutoring and simple problem-solving. Figure 1 illustrates the three key elements mentioned above.
Based on the key elements outlined, we introduce VoxEval, a SpeechQA benchmark designed to assess the knowledge understanding capabilities of SLMs with the following properties. First, VoxEval includes both questions and answers in audio format, making it, to the best of our knowledge, the first benchmark that supports end-to-end evaluation of speech-based interactions. Second, it incorporates diverse input audio conditions for the same question, allowing us to test the robustness of SLMs across various audio scenarios. Third, to tackle the challenges of evaluating subjects like mathematics and other complex topics, we identify the critical distinctions between written and spoken formats, transform written content into its spoken equivalent, and incorporate Chain-of-Thought prompts to assess the reasoning ability of SLMs in a step-by-step problem-solving setting. Table 1 presents the statistics for VoxEval, while Table 2 compares the distinctive features of VoxEval with those of other benchmarks. Our evaluation of recently developed SLMs shows that VoxEval poses a substantial challenge for current models, as they perform poorly on VoxEval, exhibit sensitivity to changes in input audio conditions, and demonstrate very limited reasoning capabilities. To summarize our contributions:
@@ 2.1 Speech Large Language Models
Benchmarks e2e # KUPs # variations math CoT
@@
Table 2: A comparison of our benchmark with other benchmarks. KUP and CoT stand for Knowledge Understanding Problem and Chain-of-Thought prompts.
@@ 2.2 SpeechLLM Evaluation Benchmarks
The evaluation of SpeechLLMs centers on their ability to model speech data, with researchers examining various aspects and levels. Initially, multiple datasets are utilized to assess the linguistic modeling capabilities of SpeechLLMs. For instance, sWUGGY Nguyen et al. (2020) is employed for lexical evaluation, sBLIMP Nguyen et al. (2020) for syntactical assessment, and sStoryCloze Hassid et al. (2024) for semantic evaluation. Additionally, the modeling of paralinguistic elements (spoken content beyond words) is crucial for SpeechLLMs, prompting the development of datasets that evaluate paralinguistic information from both token and perceptual levels Kharitonov et al. (2022); Nguyen et al. (2024). Beyond the modeling aspect, the performance of downstream applications is vital for the practical use of SpeechLLMs in real-world scenarios. Recently, numerous downstream evaluation benchmarks have emerged Sakshi et al. (2024); Wang et al. (2024a); Yang et al. (2024); Huang et al. (2024), focusing on general audio understanding, including speech, environmental sounds, and music. However, as highlighted in Section 1, these benchmarks contain all QA pairs in text format, making them suitable for evaluating S2TLLMs but not SLMs. The closest benchmark to our work is VoiceBench Chen et al. (2024), which evaluates SpeechLLMs under various audio conditions. However, VoiceBench does not support end-to-end evaluation, as it only evaluates based on textual output. Additionally, our study places a stronger emphasis on evaluating the knowledge comprehension abilities of SLMs, not only by including a broader range of topics but also by particularly focusing on their complex skills like mathematics. We also offer a detailed analysis of the performance of SLMs across various scenarios.
@@ 3.1 Data Construction
Building a speech-based knowledge understanding benchmark requires the collection of knowledge from different subjects and the construction of QA pairs about the knowledge in speech format. Given that there are numerous existing textual knowledge understanding benchmarks (see Section 2.3), it is logical to build the speech-based benchmark upon textual benchmarks. In this work, we leverage the MMLU Hendrycks et al. (2021) test set to build VoxEval. Specifically, we transform the questions from MMLU into speech using the OpenAI Text-to-Speech (TTS) API  [^2] https://platform.openai.com/docs/guides/text-to-speech. We select MMLU as the underlying dataset for the following reasons:
@@
To transform the MMLU test set into its spoken version, we need to concatenate all the text within a problem into a single sequence. In MMLU, every data is a Multiple Choice Question (MCQ) with four answer choices. We use a simple sentence to concatenate the question and answer choices. Therefore, every converted data looks like this:
@@ 3.2 Various Input Conditions
As mentioned in Section 1, it is crucial for the answers of SLMs to be robust under various input audio conditions. In this work, we define SLM robustness as the capability of an SLM to maintain consistent knowledge understanding QA accuracy when the semantic content of the input remains unchanged but the audio characteristics vary. To evaluate this robustness systematically, we incorporate MMLU questions with different input audio conditions into VoxEval. Specifically, we consider the following types of input conditions.
@@ 3.2.2 Different Speaking Styles
In real-world conversations, even a single individual may use different speaking styles depending on the situation. Therefore, SLMs need to manage various input styles. We identify two types of style variations: linguistic and paralinguistic.
@@
Models SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice Whisper + Llama-3.1-8B
@@
Alloy 0.01 4.80 20.84 12.16 37.63 55.25
Echo 0.01 5.58 20.96 12.21 37.64 54.68
Fable 0.00 1.16 20.84 11.53 36.42 53.75
Nova 0.01 3.32 20.70 12.98 36.77 54.23
Onyx 0.02 2.75 19.66 11.92 37.64 55.29
Shimmer 0.00 5.16 20.76 12.05 38.15 55.15
@@
Linguistic 0.01 4.88 20.44 11.87 36.43 55.73
Speed 0.01 5.03 19.11 10.13 34.69 54.16
Pitch 0.00 5.44 17.88 6.09 33.45 45.69
@@
Noise 0.00 3.68 19.50 10.18 36.95 55.51
Other Env Acou 0.01 4.34 20.19 10.51 37.28 55.11
@@
Models S $\rightarrow$ S S $\rightarrow$ T T $\rightarrow$ S T $\rightarrow$ T
@@
  • Speed change. In real-life situations, people often vary their speaking speed, sometimes talking faster and other times slower. To mimic these variations in speaking pace, we apply tempo changes. For each question input, we randomly select a tempo-changing rate that ranges from half the original speed to twice as fast.
@@ 3.2.3 Different Audio Qualities
  • Background noise. Sometimes the input audio may be accompanied by background noise, which can affect the clarity and intelligibility of the spoken content. We consider different kinds of noise, including Gaussian noise, colored noise, background music, and short interruption noise. For each problem, a kind of noise is randomly selected and applied to the original audio. Table 13 presents the transformation settings for all the noise types used.
  • Other environment acoustics. We consider speech recorded under various conditions, including aliasing, room impulse response, low pass filter, high pass filter, band pass filter, bitcrush, gain, clipping distortion, and seven band parameter EQ. Full explanations and transformation settings are presented in Appendix E. A random audio effect is applied to each question input.
@@
Models (MGS) SpeechGPT TWIST SPIRIT-LM Moshi GLM-4-Voice
@@
Table 7: Average accuracy (%) for various subject groups across different input audio conditions are presented. The scores are shown in the format “STEM/Social Science/Humanities/Others" subjects. MGS stands for Mean Grouped Score. The highest score in each scenario is highlighted in bold.
@@ 3.3 Handle Math Expressions and More
Mathematical problems often include numerous mathematical expressions, and current TTS systems struggle to accurately convert these expressions into speech, as they are typically trained only on textual words. To tackle this issue, we propose a two-step method for synthesizing math problems. First, we convert the written math problems into their spoken form, and then we use the TTS system to generate speech. Specifically, we create a set of few-shot examples and use GPT-4o to handle the conversion process. Our focus is on transforming Arabic numerals, units, operators, and brackets, as illustrated by the examples in Table 4. The complete prompt is provided in Figure 4 and 5.
@@
Models Direct LLM Direct LLM Direct LLM
@@
SPIRIT-LM 15.87 12.17 18.51 17.41 18.00 18.00
Moshi 15.08 16.40 12.59 13.70 15.00 18.00
GLM-4-Voice 21.96 23.02 5.19 4.81 7.00 5.00
@@
SPIRIT-LM 17.20 15.34 13.33 11.48 15.00 13.00
Moshi 16.67 16.67 12.96 13.33 14.00 12.00
GLM-4-Voice 23.02 22.49 7.78 7.41 6.00 5.00
@@ 4.1 Experimental Setups
Models. We evaluate VoxEval on five recently introduced (open-source) SLMs: SpeechGPT Zhang et al. (2023), TWIST Hassid et al. (2024), SPIRIT-LM Nguyen et al. (2024), Moshi Défossez et al. (2024), and GLM-4-Voice Zeng et al. (2024). These models represent diverse training approaches within the SLM domain. Specifically, SpeechGPT and TWIST are pre-trained using unsupervised speech data. SPIRIT-LM and GLM-4-Voice are pre-trained on interleaved speech and text data. Moshi is pre-trained on aligned speech-text data. All the models are built upon TLM checkpoints. Furthermore, while the other models process audio data through a single stream, Moshi uniquely represents audio data using multiple streams.
@@ 4.2 Results
To answer RQ1, we evaluate five SLMs on VoxEval, as shown in Table 5. Results reveal that current SLMs perform poorly, with most failing to surpass random guessing. Since VoxEval questions have four answer choices, random guessing yields an expected score of 25%. However, only GLM-4-Voice exceeds this baseline, indicating that SLMs often struggle to follow instructions and select the correct answer. This underscores the challenge of enabling SLMs to “speak" accurate responses, making VoxEval a tough benchmark. A case study of SLMs’ spoken responses is in Appendix G. Notably, SpeechGPT’s performance is near 0%. This is likely due to our approach of directly performing speech-to-speech generation without utilizing the “chain-of-modality" mode described in Zhang et al. (2023). We choose not to use the “chain-of-modality" approach because it involves converting the input into text before generating a response, which we believe would significantly increase inference latency and is not an optimal strategy for end-to-end SLMs.
@@
To answer RQ4, we present the performance of SLMs across three math subjects in VoxEval, as shown in Table 8. The selected subjects—elementary math, high school math, and college math—are intended to represent varying levels of reasoning complexity. The math-solving skills of SLMs are evaluated in two ways: direct answer, where the model performs 5-shot inference to provide the answer directly without showing intermediate reasoning, and Chain-of-Thought prompting (CoT) Wei et al. (2022), which uses annotated intermediate steps from three dev-set questions per subject as examples appended before the actual question. All the CoT examples and questions are in speech format. CoT evaluations are conducted in two settings: Cut, where the input audio is trimmed to the last 80 seconds (consistent with other experiments), and No cut, where the audio is left untrimmed to preserve longer CoT prompts. Two metrics are used for evaluation: string matching, the standard final answer extraction method, and LLM, where GPT-4o Hurst et al. (2024) is used to extract answers to complement string matching, as models often include lengthy reasoning steps that hinder accurate string-based extraction when applying CoT. Our observations include:
  • Existing SLMs perform similarly across different levels of math questions. Interestingly, in some cases, these models occasionally achieve better results on more complex problems than on simpler ones. Among them, GLM-4-Voice stands out as the only SLM consistently excelling in simple (elementary) math. This counterintuitive fact suggests that most SLMs are unlikely to possess basic math reasoning abilities.
@@ 6 Limitations
The limitations of this work come from the scope and the data construction methodology. First, the VoxEval benchmark focuses exclusively on assessing the robustness of SLMs in terms of knowledge understanding, leaving out other critical aspects such as fairness, toxicity, hallucination, etc. These issues are particularly relevant in generative scenarios, where SLMs may exhibit biases, generate harmful content, or produce inaccurate information. However, these aspects fall outside the scope of VoxEval and require separate evaluation frameworks tailored to address them comprehensively. Second, the benchmark’s reliance on synthetic speech generation through TTS systems may not fully capture the complexity and variability of natural human speech, including emotional expressions, regional dialects, and spontaneous speech patterns that are difficult to replicate artificially.
Acknowledgments
The research presented in this paper was partially funded by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23).
Appendix A Accuracy of the Answer Extraction Pipeline
Models GLM-4-Voice Moshi SPIRIT-LM
WER 0.01 0.16 0.08
CER 0.01 0.10 0.04
Total Word Errors 3 12 23
Total Character Errors 13 22 64
Total Words in GT 300 78 308
Total Characters in GT 1759 401 1767
Table 9: ASR process evaluation results. GT represents Ground Truth.
To ensure the reliability of our automated pipeline (ASR + string matching) for extracting final answers from SLMs’ speech output, it is essential to assess its accuracy. This section provides the evaluation results, focusing on two key questions. 1) How accurate is the ASR process in transcribing the SLMs’ spoken output into text? 2) Is the final answer extracted correctly?
To answer the above two questions, we randomly select ten questions from the VoxEval benchmark and conduct spoken QA using three top-performing SLMs (i.e., GLM-4-Voice, Moshi, and SPIRIT-LM). This process results in 30 SLM-generated responses. For the first question, we use Whisper-large-v3 to transcribe these 30 responses and also manually transcribe them to serve as ground truth. Finally, we calculate the Word Error Rate (WER) and Character Error Rate (CER) by comparing the ASR transcriptions with the manual ground truth. As shown in Table 9, all three models have very low WER and CER scores. While Moshi exhibits relatively higher error rates, this can be attributed to its characteristically shorter responses compared to the other models. Since WER and CER are normalized by the ground truth length, direct comparison may not fully reflect relative performance. Nevertheless, these results indicate minimal ASR transcription errors across all models. For the second question, using the same 30 model responses, we apply our automatic evaluation pipeline to extract final answers and compare these against manually extracted ground truth answers obtained through direct listening. The accuracy achieved is 29/30. The single error occurred with a SPIRIT-LM response characterized by rapid and unclear speech, resulting in ASR failure. However, this error is attributable to the model’s failure to provide a clear, intelligible answer rather than a systematic pipeline limitation. To summarize, these experiments demonstrate the robustness and reliability of our automatic evaluation pipeline, with high accuracy in both ASR transcription and the answer extraction task.
Appendix B Comparison between Different Speech Interaction Models
Models Accuracy
GLM-4-Voice 0.472
Whisper + Llama-3.1 0.606
Whisper + Qwen2 0.517
gpt-4o-audio-preview 0.718
Table 10: Performance comparison between different speech interaction models.
To better understand how well existing end-to-end SLMs are, we compare their performance with other types of speech interaction models. Specifically, we examine two alternative model categories:
  1. 1.
    Open-source cascaded models: These models follow a pipeline structure that combines ASR with a TLM, so they are not considered end-to-end SLMs. This comparison helps determine whether cascaded models or SLMs perform better. For this category, we use two combinations: Whisper-large-v3 + Llama-3.1-8B-Instruct and Whisper-large-v3 + Qwen2-7B-Instruct.
  2. 2.
    Closed-source API models: These are speech interaction models whose checkpoints are not publicly available and can only be accessed through their official APIs. This comparison aims to identify performance gaps between open-source and closed-source models. For this category, we use gpt-4o-audio-preview.
We include only GLM-4-Voice for SLMs, as it performs the best in our main experiments. Due to budget constraints, we choose three subjects from each of the four subject groups (STEM, social science, humanities, others) of VoxEval, resulting in 12 subjects. For each subject, we keep the first 100 QA pairs for this evaluation.
We observe two key findings from the results shown in Table 10. First, cascaded models outperform SLMs, highlighting unique challenges associated with developing end-to-end models. Second, the closed-source API model achieves the highest performance and demonstrates a notable gap compared to SLMs, suggesting the potential performance ceiling for SLMs is far from being reached.
Appendix C Further Analysis of the In-context Learning Capabilities of SLMs with and without CoT
Elementary Math High School Math College Math
Models Non-CoT CoT Non-CoT CoT Non-CoT CoT
0-shot 0.07 0.07 0.06 0.06 0.02 0.02
1-shot 0.25 0.28 0.20 0.04 0.25 0.08
2-shot 0.24 0.31 0.21 0.05 0.27 0.11
3-shot 0.26 0.28 0.26 0.08 0.27 0.08
4-shot 0.25 - 0.27 - 0.25 -
5-shot 0.23 - 0.26 - 0.31 -
Table 11: In-context learning results of GLM-4-Voice with and without CoT. Non-CoT represents in-context learning without CoT (direct answer).
To gain deeper insights into whether current SLMs benefit from in-context learning, we conduct additional experiments on GLM-4-Voice. Specifically, we evaluate three mathematical domains in VoxEval using varying numbers of in-context learning examples, assessing GLM-4-Voice performance both with and without Chain-of-Thought (CoT) prompts. We use the three subject data in Appendix B. Our analysis of the results presented in Table 11 reveals three key findings.
  • In-context examples improve performance for both CoT and non-CoT approaches. The 1-shot configurations consistently outperform 0-shot baselines across both CoT and non-CoT settings, demonstrating that SLMs can effectively leverage contextual examples to enhance their mathematical reasoning capabilities.
  • CoT effectiveness varies inversely with problem complexity. While CoT prompting consistently outperforms non-CoT approaches in elementary mathematics, this advantage diminishes for more advanced topics (high school and college mathematics). This pattern suggests that GLM-4-Voice can effectively utilize reasoning paths provided in CoT prompts for simpler problems, indicating the presence of reasoning capabilities in SLMs. However, these benefits do not extend to more complex mathematical domains.
  • GLM-4-Voice exhibits a trade-off between instruction-following and reasoning abilities. Notably, 1-shot CoT underperforms 1-shot non-CoT on more challenging mathematical problems. This phenomenon suggests that in non-CoT settings, SLMs may rely on simple guessing rather than explicit reasoning. Conversely, when CoT prompts guide the model toward explicit reasoning on problems beyond its reasoning capacity, the model may become overly focused on the reasoning process at the expense of proper answer formatting and selection.
Appendix D Statistical Test Results across Different Input Audio Conditions
Models $\chi^{2}$(10) p-value
GLM-4-Voice 121.53 2.47e-21
Moshi 180.54 1.81e-33
SPIRIT-LM 41.09 1.09e-05
TWIST 254.76 5.40e-49
SpeechGPT 7.97 0.63
Table 12: Friedman Test results for the five SLMs evaluated in this paper.
(a) GLM-4-Voice
(b) Moshi
(c) SPIRIT-LM
(d) TWIST
Figure 3: Wilcoxon signed-rank test results for the four SLMs showing statistically significant differences in the Friedman test.
To better understand how various input audio conditions influence the performance of different SLMs, we conduct statistical analyses on the models’ responses across the 11 distinct audio conditions (see Table 5) for each SLM. Specifically, we collect the 56 subject accuracy scores of each SLM under each audio condition and apply the Friedman Test to analyze all conditions, followed by Wilcoxon Signed-Rank Tests for pairwise comparisons between conditions. A Type I error $\alpha$ of 5% is used in the statistical analyses.
The Friedman Test results, summarized in Table 12, reveal that input audio conditions significantly affect the performance of all SLMs except SpeechGPT, which consistently exhibits near-zero accuracy across all conditions. To further investigate these effects, we conduct pairwise Wilcoxon Signed-Rank Tests with Bonferroni correction on the remaining four models. The corrected significance level is calculated as:
$\alpha_{corrected}=\frac{\alpha}{n}=\frac{0.05}{55}=0.000909,$ (1)
where
$n=\binom{11}{2}=55$ (2)
represents the total number of pairwise comparisons among the 11 audio conditions. The resulting p-values are visualized as heatmaps in Figure 3, with statistically significant values highlighted in bold. The heatmaps demonstrate that the impact of input audio conditions varies across different SLMs. Notably, GLM-4-Voice, Moshi, and SPIRIT-LM exhibit significant performance differences under the "Speed" and "Pitch" conditions, suggesting that these models are particularly sensitive to variations in speaking rate and pitch. This finding underscores the importance of robust handling of diverse user speech patterns in SLM design.
3:  for each input audio condition setting $S_{i}$ do
5:     Initialize an empty list $D_{i}=[]$ to store difference scores for $S_{i}$.
6:     for each subject $n\in\{1,2,\dots,N\}$ do
8:        Compute the maximum absolute score difference:
9:        Append $d_{n}$ to $D_{i}$: $D_{i}\leftarrow D_{i}\cup\{d_{n}\}$.
10:     end for
11:     Use $D_{i}=\{d_{1},d_{2},\dots,d_{56}\}$ to generate a box plot for $S_{i}$.
12:  end for
Appendix E Explanation for Various Speech Recording Conditions